A 150 lb., 62-year-old man had a chronic productive cough, exertional dyspnea, mild cyanosis, and marked slowing of forced expiration. Calculate the tidal volume (TV) for this person before and after the bronchodilator therapy.

His pulmonary function and laboratory tests follow:

Frequency 16 breaths/min
Alveolar ventilation 4.2 L/min
Vital capacity (VC) 2.2 L
Functional residual capacity (FRC) 4.0 L
Total lung capacity (TLC) 5.2 L
Maximum inspiratory flow rate 250 L/min
Maximum expiratory flow rate 20 L/min
PaO2 62 mm Hg
PaCO2 39 mm Hg

Pulmonary function tests after bronchodilator therapy:

Frequency 16 breaths/min
Alveolar ventilation 4.35 L/min
VC 2.4 L
FRC 4.0 L
TLC 5.2 L
Maximum inspiratory flow rate 250 L/min
Maximum expiratory flow rate 23 L/min
PaO2 62 mm Hg
PaCO2

38 mm Hgc
Answer #1

Tidal volume is the lung volume representing the normal volume of air displaced between normal inhalation and exhalation when extra effort is not applied.

Formula :

Vital capacity = Inspiratory reserve volume + Tidal volume + Expiratory resrve volume

Vital capacity before bronchodilator therapy is 2.2 L

IRV before bronchodilator therapy is 250 L/min

ERV before bronchodilator therapy is 20 L/min

Applying this to the formula , 2.2L = 250 L/min + TV + 20L/min

2.2= 270 +TV

TV = 270-2.2 = 268.2 L/min

Vital capacity = Inspiratory reserve volume + Tidal volume + Expiratory resrve volume

Vital capacity after bronchodilator therapy is 2.4 L

IRV after bronchodilator therapy is 250 L/min

ERV after bronchodilator therapy is 23 L/min

Applying this to the formula , 2.4 L = 250 L/min + TV + 23 L/min

2.4= 273 +TV

TV = 273 -2.4 = 271.4 L/min